Say true or false and justify your answer: 
 2^3 > 5^2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us consider LHS = 2^3
Expansion of 2^3= 2 × 2 × 2=8
Now, consider RHS = 5^2
Expansion of 5^2= 5 × 5 = 25
By comparing LHS and RHS, LHS < RHS
2^3< 5^2
Hence, the given statement is false.

Thanks